# Adaptive Collaboration Strategy for LLMs in Medical Decision Making

### v1

@@ 3.4 Decision Making
MedVidQA v + t Question + Answer A/B/C/D 155 First Aids, Medical Emergency, and Medical Education Questions
4 Experiments and Results
In this section, we evaluate our framework and baseline models using state-of-the-art LLMs across multiple medical benchmarks in Solo, Group, and Adaptive settings. Our experiments and ablation studies highlight the framework’s dynamic performance, demonstrating robustness and efficiency by modulating agent numbers and temperature. Results also show a beneficial convergence of agent opinions in collaborative settings.
Figure 3: Our adaptive setting outperforms Solo and Group settings on multiple medical benchmarks across LLMs. Note that with GPT-4 (V), Our adaptive setting performs the best in all medical benchmarks except for the DDXPlus.
Table 3: The best result of each LLM with different agent settings evaluated on medical datasets. t: text, i: image, v: video.
GPT-3.5 GPT-4(V) Gemini-Pro(Vision)
Dataset Modality Solo Group Ours Solo Group Ours Solo Group Ours
MedQA t-only 64.0 63.3 80.0 88.0 83.7 95.0 52.5 56.3 74.8
MedMCQA t-only 73.0 60.0 66.0 85.0 84.0 100.0 66.7 60.0 63.0
PubMedQA t-only 73.4 58.2 80.0 63.0 77.1 94.0 67.9 82.5 72.0
DDxPlus t-only 68.8 77.8 74.0 83.8 73.3 80.0 65.1 77.3 78.9
PMC-VQA i + t - - - 55.0 44.4 57.6 48.9 40.5 44.0
Path-VQA i + t - - - 64.9 67.1 68.8 66.7 67.1 86.0
MedVidQA v + t - - - - - - 92.3 77.8 86.2
4.1 Setup
To verify the effectiveness of our framework, we conduct comprehensive experiments with baseline methods on seven datasets including MedQA, MedMCQA, PubMedQA, DDXPlus, Path-VQA, PMC-VQA and MedVidQA. Table 2 summarizes the data statistics. More information about the datasets and tasks is presented in Appendix A.
In all the experiments, we evaluate the performance of agents driven by GPT-3.5 [51, 4], GPT-4 (v) [50], and Gemini-pro (vision) [66] across various tasks. For all the quantitative experiments in this section, we compare three settings: (1) Solo: Using a single LLM agent in the decision-making state. (2) Group: Implementing a group of agents to collaborate during the decision-making process. (3) Adaptive: Our proposed method MDAgents, adaptively constructs the inference structure from PCP to MDT and ICT.
Medical Question Answering
With MedQA [29], MedMCQA [52] and PubMedQA [30], we focus on question answering through text, involving both literature-based and conceptual medical knowledge questions. Specifically, PubMedQA tasks models to answer questions using abstracts from PubMed, requiring synthesis of biomedical information. MedQA and MedMCQA test model’s ability to understand and respond to multiple-choice questions derived from medical educational materials and examinations. These tasks evaluate the AI’s medical reasoning and comprehension capabilities.
Diagnostic Reasoning
DDxPlus [65] involves clinical vignettes that require differential diagnosis, closely mimicking the diagnostic process of physicians. The task tests the model’s ability to reason through symptoms and clinical data to suggest possible medical conditions, thus evaluating the AI’s diagnostic reasoning abilities similar to a clinical setting.
Medical Visual Interpretation
Path-VQA [25], PMC-VQA [86] and MedVidQA [23] datasets challenge models to interpret medical images and videos, requiring integration of visual data with clinical knowledge. In detail, PathVQA focuses on answering questions based on pathology images, testing AI’s capability to interpret complex visual information from medical images. PMC-VQA evaluates AI’s proficiency in deriving answers from both text and images found in scientific publications, requiring a multifaceted understanding of visual and textual content. Finally, MedVidQA extends to video-based content, where AI models need to process information from medical procedure videos, combining visual cues, and narrative understanding to answer related questions.
Baseline Methods
  • Solo: The baseline methods considered for the Solo setting include the followings: Zero-shot [33] directly incorporates a prompt to facilitate inference, while Few-shot [4] involves a small number of examples. Few-shot CoT [76] integrates rationales before deducing the answer. Few-shot CoT-SC [74] builds upon Few-shot CoT by sampling multiple chains to yield the majority answer. Ensemble Refinement (ER) [60] is a prompting strategy that conditions model responses on multiple reasoning paths to bolster the reasoning capabilities of LLMs. Medprompt [49] is a composition of several prompting strategies that enhances the performance of LLMs and achieves state-of-the-art results on multiple benchmark datasets, including medical and non-medical domains.
  • Group: We tested five group-based decision-making methods: Voting [75], Multidisciplinary Collaboration (MC) [64], Reconcile [6], AutoGen [80], and DyLAN [41]. Autogen was based on four agents, with one User Agent, one Clinician, one Medical Expert, and one Moderator. Each agent was given one round of respond  [^1] https://microsoft.github.io/autogen/docs/notebooks/agentchat_groupchat. DyLAN setup followed the base implementations of four agents with no specific roles and four maximum rounds of interaction.  [^2] https://github.com/SALT-NLP/DyLAN/tree/main
4.2 Main Results
In Table 3, we report the classification accuracy on MedQA, MedMCQA, PubMedQA, DDXPlus, Path-VQA, PMC-VQA and MedVidQA dataset. We compare our method (Adaptive) with several baselines in both Solo and Group settings.
Adaptive method outperforms Solo and Group settings.
As depicted in Figure 3 and Table 3, MDAgents outperforms baseline methods within both Solo and Group settings, showing best performance in 5 out of 7 medical benchmarks tested. This reveals the effectiveness of adaptive strategies integrated within our system, particularly when navigating through the text-only (e.g., MedQA where it outperformed Solo by up to 22.3% and Group by up to 18.5%) and text-image datasets (e.g., Path-VQA and PMC-VQA). Our approach not only comprehends textual information with high precision but also adeptly synthesizes visual data, a pivotal capability in medical diagnostic evaluations.
Why Adaptive Decision Making Framework Works Well
Figure 4: Experiment with MedQA dataset with N=10 randomly sampled questions. (a) The average plot for the probability of moderator provides the complexity level with the best / middle / worst choice for each problem. (b-d) The distribution of accuracy level of low / moderate / high solver. We solved each question with a specific complexity level 10 times.
We hypothesize that an LLM, functioning as a classifier or a generalist doctor, will select the optimal complexity level for each MDM problem. This hypothesis is supported by Figure 4, which illustrates that the model appropriately matches the complexity levels—low, moderate, and high—of the given problem.
[…]
Solo vs. Group setting.
In Table 4 and 5 in Appendix, we show performance comparison within Solo and Group settings. When individual agents demonstrate competent performance, such as in MedQA, MedVidQA, and DDxPlus, the precision of solo decision-making often outperforms group methods, attributing to a reduced cognitive load and avoidance of groupthink, a state where the desire for harmony in a decision-making group results in irrational or dysfunctional outcomes [46]. On the other hand, as the complexity of tasks increases, group settings can suffer from polythink, where excessive perspectives lead to confusion and collective judgment errors, reflecting the challenges seen in self-debugging processes under complex conditions [24]. The group strategies, while robust in information-rich environments, appear vulnerable to compounded inaccuracies during collaborative reasoning, especially in complex problem domains [24]. Our adaptive approach avoids these shortcomings by flexibly choosing the reasoning strategies, effectively merging the benefits of both individual and collective decision-making. It constantly improves its performance by discussion and review, which ensures consistently high accuracy across various medical datasets.
Convergence Trends in Consensus Dynamics.
As depicted in Figure 5, the trend towards consensus among MDAgents is evident across different data modalities. The text+video modality demonstrates a rapid convergence, potentially reflecting the agents’ efficient processing of combined textual and visual cues. On the other hand, the text+image and text-only modalities display a more gradual decline in entropy, indicating a progressive narrowing of interpretative diversity among the agents. Despite the differing rates and initial conditions, all modalities exhibit a convergence of agent opinions over time. This uniformity in reaching consensus highlights the MDAgents’ capability to integrate and reconcile information, irrespective of the data format. The entropy H here serves as an indicator of consensus progression, quantified as:
$H=-\sum_{i=1}^{M}p(x_{i})\log_{2}p(x_{i})$ (4.1)
where M is the total number of unique answers, $x_{i}$ represents a unique answer, and $p(x_{i})$ is the probability of occurring among all answers.
4.3 Ablation Studies
Impact of Number of Agents in Group Setting.
In Figure 6, we examine the effects of varying the number of agents in a collaborative Group setting. The result shows that a higher number of agents does not leads to better performance; rather, our Adaptive method, which intelligently calibrates the number of collaborating agents, achieves optimal performance with fewer agents (N=3), as shown by its peak accuracy of 83.5%. This not only indicates efficiency in decision-making but also computational and economic benefits, considering the reduced number of API calls needed, especially when contrasted with the Solo and Group settings.
With regards to computational efficiency and cost, the Solo setting (5-shot CoT-SC) resulted in a 6.0 and Group setting (Delphi with N=5) resulted in a 20.3 API calls, suggesting a high computational cost without a corresponding increase in accuracy. On the other hand, our Adaptive method exhibits a more economical use of resources, demonstrated by fewer API calls (13.6 with five agents) while maintaining high accuracy, a critical factor in deploying scalable and cost-effective medical AI solutions.
Robustness of MDAgents with different parameters.
In Figure 6, we highlight the robustness of different LLM settings under low (T=0.3) and high (T=1.2) temperature conditions. Our Adaptive setting shows remarkable resilience to changes in temperature, with performance improving under higher temperatures. This suggests that our model can utilize the creative and diverse outputs generated at higher temperatures to enhance decision-making, a property that is not as pronounced in the Solo and Group settings. Such robustness is particularly valuable in real-world medical applications, where varying degrees of uncertainty and ambiguity in data are the norm.
Impact of Complexity Selection.
Figure 7 shows the impact of query complexity on model accuracy. Our adaptive method significantly outperforms the static complexity assignments (Low, Moderate, High) across benchmarks with different modality inputs, highlighting the necessity for dynamic complexity assessment in MDM. In text-only queries, the Low setting presents a close second to the Adaptive method, underscoring that for less complex queries, simpler models can perform adequately. However, the real testament to the Adaptive model’s superiority is within the text+image and text+video modalities, where it achieves 83.3% and 92.9% accuracy, respectively, far exceeding the Low and other static models. This emphasizes the model’s agility in handling the compounded difficulty of multi-modal medical data.
5 Conclusion
This paper introduces MDAgents, a framework designed to enhance the utility of LLMs in complex medical decision-making by dynamically structuring effective collaboration models. To reflect the nuanced consultation aspects in clinical settings, MDAgents adaptively assigns LLMs either to roles independently or within groups, depending on the task’s complexity. This emulation of real-world medical decision processes has been comprehensively evaluated, with MDAgents outperforming previous solo and group methods in 5 out of 7 medical benchmarks. The case study illustrates the practical efficacy and collaborative dynamics of our proposed framework, providing insights into how differing expert opinions are synthesized to reach a more accurate diagnosis. This is evidenced by our agents’ ability to converge on the correct diagnosis despite initially divergent perspectives. Ablation studies further elucidate the individual contributions of agents and strategies within the system, revealing the critical components and interactions that drive the framework’s success. By harnessing the strength of multi-modal reasoning and fostering a collaborative process among LLM agents, our framework opens up new possibilities for enhancing LLM-assisted medical diagnosis systems, pushing the boundaries of automated clinical reasoning.
Limitations and Future Works
Despite the successes of our framework in showing promising performance in medical decision-making tasks, we recognize several limitations that open pathways for future research.
Medical Focused Foundation Models.
An essential enhancement would be to incorporate the foundation models specifically trained on medical data, like Med-PaLM 2 [60]. These models will serve as a more accurate and up-to-date knowledge base, tailored to understand and generate medically accurate content with reduced error margins. By grounding the LLMs in the medical domain, we aim to diminish hallucinations and enrich the model’s expertise.
Patient-Centered Diagnosis.
A primary limitation lies in the fact that our current framework operates within the confines of multi-choice question answering and does not account for patient-physician interactions. Real-world medical diagnostics are deeply rooted in patient-centered approaches, where differential diagnosis is informed by a continuous exchange between the patient’s narrative and the physician’s expertise. To bridge this gap, future iterations of our framework will aim to incorporate a more interactive system that not only assists physicians but also engages directly with patients. By embedding real physicians and patients within the feedback loop, the system can be trained to become more nuanced and patient-centric, thereby enhancing the quality and personalization of medical decision support.
Updating Medical Knowledge.
Another limitation is the static nature of the LLMs’ medical knowledge. As medicine is a rapidly evolving field, the information within the LLM must be continually updated to reflect the latest medical insights and findings. To address this, we propose the integration of a medical Retrieval-Augmented Generation (RAG) technique. This addition would allow the framework to be pulled from up-to-date medical literature, ensuring that the model’s responses and decisions are informed by the most current data and evidence-based practices.
Acknowledgments and Disclosure of Funding
We thank Yoon Kim at MIT, Vivek Natarajan at Google, WonJin Yoon at Harvard Medical School, Seonghwan Bae at Sacheon Public Health Center, Hyeonhoon Lee and Hui Dong Lim at Seoul National University Hospital for their revisions, feedback, and support.
Appendix A Dataset Information
@@
    Sample Question: "A 47-year-old female undergoes a thyroidectomy for treatment of Graves’ disease. Post-operatively, she reports a hoarse voice and difficulty speaking. You suspect that this is likely a complication of her recent surgery. What is the embryologic origin of the damaged nerve that is most likely causing this patient’s hoarseness?”
    Options: "A: 1st pharyngeal arch, B: 2nd pharyngeal arch, C: 3rd pharyngeal arch, D: 4th pharyngeal arch, E: 6th pharyngeal arch"
@@
    MedMCQA. Similarly, the MedMCQA is derived from real-world medical entrance exams and includes approximately 194K multiple-choice questions across 21 medical subjects. We use 6.1K samples from the test set. Given the similarity to MedQA in terms of format and content, this dataset is also classified as having low complexity.
    Sample Question: “Retraction of mandible is achieved by: ”
    Options: “Lateral pterygoid", “Temporalis", "Medial pterygoid", and "Masseter"
@@
    PubMedQA. PubMedQA is a QA dataset based on biomedical research [30]. It requires yes/no/maybe answers to questions grounded in PubMed abstract. The dataset comprises entries each containing a question, a relevant abstract minus the conclusion, and a ground truth label. We used 500 samples for testing. Given its binary choice format, we consider the complexity of this dataset to be low.
    Context: "The use of a balloon expandable stent valve includes balloon predilatation of the aortic stenosis before valve deployment. The aim of the study was to see whether or not balloon predilatation is necessary in transcatheter aortic valve replacement (TAVI). Sixty consecutive TAVI patients were randomized to the standard procedure or to a protocol where balloon predilatation was omitted. There were no significant differences between the groups regarding early hemodynamic results or complication rates."
  4. 4.
@@
    Evidences: [’e66’, ’insp_siffla’, ’j45’, ’posttus_emesis’, ’trav1_@_N’, ’vaccination’]"
@@
Appendix B Prompt Templates
B.1 A single agent setting
@@
B.2 Multi-agent setting
You are a medical expert who conducts initial assessment and your job is to decide the difficulty/complexity of the medical query. Now, given the medical query as below, you need to decide the difficulty/complexity of it: {{question}} Please indicate the difficulty/complexity of the medical query among below options: 1) low: a single medical agent can output an answer. 2) moderate: number of medical experts with different expertise should dicuss and make final decision. 3) high: multiple teams of clinicians from different departments need to collaborate with each other to make final decision. **Answer:**( Complexity check prompt
@@ B.2 Multi-agent setting
You are a final medical decision maker who reviews all opinions from different medical experts and make final decision. Given the {{inputs}}, please review the {{inputs}} and make the final answer to the question by {{decision_methods}}. **Answer:**( Final decision prompt
Table 4: Accuracy (%) on Medical benchmarks with Solo setting. CoT refers to Chain-of-Thought, SC refers to Self-Consistency, ER refers to Ensemble Refinement prompting. Bold number represents the best performance for each benchmark and model.
Model Prompting Dataset
MedQA MedMCQA PubMedQA DDxPlus Path-VQA PMC-VQA MedVidQA
GPT-3.5 Zero-shot 50.4 65.0 71.8 53.0 - - -
Few-shot 55.9 57.0 67.6 46.0 - - -
+ CoT 60.7 52.0 71.4 43.8 - - -
+ CoT-SC 64.0 57.0 73.4 28.2 - - -
ER 58.0 58.0 13.0 68.8 - - -
Medprompt 56.0 73.0 70.0 55.0 - - -
GPT-4(V) Zero-shot 74.0 78.0 60.0 69.0 56.0 52.0 -
Few-shot 75.5 78.0 60.0 69.0 55.0 53.0 -
+ CoT 83.0 81.0 63.0 83.8 62.2 49.0 -
+ CoT-SC 88.0 85.0 50.0 55.6 64.9 55.0 -
ER 88.0 82.0 39.0 77.6 64.9 52.0 -
Medprompt 80.0 82.0 37.0 75.5 62.0 49.5 -
Gemini-Pro(Vision) Zero-shot 48.0 57.1 51.0 48.8 63.8 41.2 40.0
Few-shot 45.5 56.7 55.0 36.6 62.1 44.4 55.0
+ CoT 48.5 66.7 66.7 65.1 50.0 41.3 30.4
+ CoT-SC 50.0 61.3 60.0 60.4 63.3 48.9 88.9
ER 52.5 63.5 67.9 34.8 66.7 45.7 92.3
Medprompt 49.5 62.2 50.0 58.3 57.4 40.0 73.3

### v2

@@ Appendix A Dataset Information
MedVidQA T V Question + Answer A/B/C/D 155 First Aids, Medical Emergency, and Medical Education Questions
@@
    Sample Question: “A 47-year-old female undergoes a thyroidectomy for treatment of Graves’ disease. Post-operatively, she reports a hoarse voice and difficulty speaking. You suspect that this is likely a complication of her recent surgery. What is the embryologic origin of the damaged nerve that is most likely causing this patient’s hoarseness?”
    Options: A: 1st pharyngeal arch, B: 2nd pharyngeal arch, C: 3rd pharyngeal arch, D: 4th pharyngeal arch, E: 6th pharyngeal arch
@@
    PubMedQA. PubMedQA is a QA dataset based on biomedical research [35]. It requires yes/no/maybe answers to questions grounded in PubMed abstract. The dataset comprises entries each containing a question, a relevant abstract minus the conclusion, and a ground truth label. We used 50 samples for testing. Given its binary choice format, we consider the complexity of this dataset to be low.
    Context: “The use of a balloon expandable stent valve includes balloon predilatation of the aortic stenosis before valve deployment. The aim of the study was to see whether or not balloon predilatation is necessary in transcatheter aortic valve replacement (TAVI). Sixty consecutive TAVI patients were randomized to the standard procedure or to a protocol where balloon predilatation was omitted. There were no significant differences between the groups regarding early hemodynamic results or complication rates."
@@
    Evidences: [‘e66’, ‘insp_siffla’, ‘j45’, ‘posttus_emesis’, ‘trav1_@_N’, ‘vaccination’]
@@
  7. 7.
    SymCat. SymCat is a synthetic dataset which includes 5 million symptom-condition samples, covering 801 distinct conditions each with 376 potential symptoms dataset [2].
    Sample Patient Information: Age: 42, Gender: F, Race: White, Ethnicity: Nonhispanic
    Sample Patient Symptoms: Pain:cramping:Abdomen:Lower abdomen::::Worsening after meals::32; Altered stool:Lumpy:::::::Often:44; Flatulence:::::::::30; Bloating:::::::::41
    Options: "A": "Asthma", "B": "IBS (Constipation type)", "C": "Viral meningitis (Varicella zoster virus)", "D": "Bacterial (Gastro)enteritis (Yersinia infection most likely)"
  8. 8.
    JAMA. JAMA includes 1524 clinical cases collected from the JAMA Network Clinical Challenge archive, which are summaries of actual challenging clinical cases. Each sample is framed as a question, with a long case description and four options [9].
    Sample Case: A 62-year-old woman undergoing peritoneal dialysis (PD) for kidney failure due to IgA nephropathy presented to the PD clinic with a 1-day history of severe abdominal pain and cloudy PD fluid. Seven days prior, she inadvertently broke aseptic technique when tightening a leaking connection of her PD catheter tubing. On presentation, she was afebrile and had normal vital signs. Physical examination revealed diffuse abdominal tenderness. Cloudy fluid that was drained from her PD catheter was sent for laboratory analysis (Table 1).Await peritoneal dialysis fluid culture results before starting intraperitoneal antibiotics.
    Sample Question: What Would You Do Next?
    "A": "Administer empirical broad-spectrum intraperitoneal antibiotics", "B": "Administer empirical broad-spectrum intravenous antibiotics", "C": "Await peritoneal dialysis fluid culture results before starting intraperitoneal antibiotics", "D": "Send blood cultures"
  9. 9.
    Medbullets. Medbullets comprises 308 USMLE Step 2/3 style questions collected from open-access tweets on X (formerly Twitter) since April 2022. The difficulty is comparable to that of Step 2/3 exams, which emulate common clinical scenarios [9].
    Sample Question: A 2-week-old boy is evaluated by his pediatrician for abnormal feet. The patient was born at 39 weeks via vaginal delivery to a G1P1 29-year-old woman. The patient has been breastfeeding and producing 5 stools/day. He is otherwise healthy. His temperature is 99.50̆0b0F (37.50̆0b0C), blood pressure is 60/38 mmHg, pulse is 150/min, respirations are 24/min, and oxygen saturation is 98% on room air. A cardiopulmonary exam is notable for a benign flow murmur. A musculoskeletal exam reveals the findings shown in Figure A. Which of the following is the most appropriate next step in management?
    Options: "A": "Botulinum toxin injections", "B": "Reassurance and reassessment in 1 month", "C": "Serial casting", "D": "Surgical pinning",
  10. 10.
    MIMIC-CXR-VQA. MIMIC-CXR is a large-scale visual question-answering dataset of 377,110 chest radiographs. It was obtained from 227,827 imaging studies sourced from the BIDMC between 2011-2016. It includes patient identifiers which can be linked to MIMIC-IV [3].
    Sample Question: Are there any abnormalities in the upper mediastinum?
    Image: p11/p11218589/s59138139/7a1e4762-c176bd78-6281fe5c-6b0c9734-e9a4c8f1.jpg
Appendix B Entropy Calculation for Consensus Dynamics
The entropy H serves as an indicator of consensus progression among the agents. It is quantified as:
$H=-\sum_{i=1}^{M}p(x_{i})\log_{2}p(x_{i})$ (B.1)
where M is the total number of unique answers, $x_{i}$ represents a unique answer, and $p(x_{i})$ is the probability of occurring among all answers. This calculation helps to measure the degree of agreement among the agents over time, with lower entropy indicating higher consensus. The trends in entropy across different data modalities provide insights into how quickly and effectively MDAgents can reach a unified decision.
Appendix C Prompt Templates
C.1 A single agent setting
@@
C.2 Multi-agent setting
You are a medical expert who conducts initial assessment and your job is to decide the difficulty/complexity of the medical query. Now, given the medical query as below, you need to decide the difficulty/complexity of it: {{question}} Please indicate the difficulty/complexity of the medical query among below options: 1) low: a PCP or general physician can answer this simple medical knowledge checking question without relying heavily on consulting other specialists. 2) moderate: a PCP or general physician can answer this question in consultation with other specialist in a team. 3) high: Team of multi-departmental specialists can answer to the question which requires specialists consulting to another department (requires a lot of team effort to treat the case). **Answer:**( Complexity check prompt
@@ C.2 Multi-agent setting
You are a final medical decision maker who reviews all opinions from different medical experts and makes the final decision. Given the {{inputs}}, please review the {{inputs}} and make the final answer to the question by {{decision_methods}} (e.g., majority voting, ensemble refinement). **Answer:**( Final decision prompt
Given the {{agent answers}}, please complete the following steps: 1. Take careful and comprehensive consideration of the provided reports. 2. Extract key knowledge from the reports. 3. Derive a comprehensive and summarized analysis based on the extracted knowledge. 4. Generate a refined and synthesized report based on your analysis. Report Generation
Appendix D Additional Results
This section presents additional experimental results and analyses that provide further insights into the performance and characteristics of our MDAgents framework.
D.1 Accuracy on entire MedQA 5-options Dataset
To provide a comprehensive evaluation of our approach, we conducted experiments on the entire MedQA 5-options dataset using GPT-4o mini. This expands upon the subsampled experiments presented in the main experiments in Table 2. Below Table shows the accuracy results for various methods.
Table 5: Accuracy (%) on entire MedQA 5-options dataset with GPT-4o mini
. Category Method Accuracy (%) Single-agent Zero-shot 71.5 3-shot 72.3 + CoT 76.6 + CoT-SC 77.2 Multi-agent (Multi-model) Majority Voting 76.3 Weighted Voting 79.1 Borda Count 76.1 Multi-agent (Single-model) Reconcile 80.2 Adaptive MDAgents (Ours) 83.6
These results demonstrate that our MDAgents approach outperforms both single-agent and other multi-agent methods across the full dataset, achieving an accuracy of 83.6%. This underscores the effectiveness of our framework in handling diverse medical questions at scale.
D.2 Estimated Costs for Full Test Set Experiments
To provide transparency and aid in reproducibility, we estimated the costs associated with running experiments on the entire test sets using GPT-4 (Vision). Below Table presents these cost estimates in USD for various datasets and methods.
Table 6: Estimated costs for experimenting with entire test sets with GPT-4 (Vision) (in USD)
Method MedQA PubMedQA Path-VQA PMC-VQA DDXPlus SymCat JAMA MedBullets MIMIC-CXR Total Cost
CoT 55.24 13.16 3,028.54 27,134.00 16,461.90 10,593.99 134.55 61.23 1,388.70 58,871.29
Ours 172.43 41.36 9,369.45 82,194.34 44,814.97 31,176.05 367.13 161.70 4,406.90 172,704.33
While our approach incurs higher costs due to its multi-agent nature, the significant performance improvements justify this increased computational expense for critical medical decision-making tasks.
D.3 Impact of Knowledge Enhancement with RAG
Method Accuracy (%)
MDAgents (baseline) 71.8
+ MedRAG 75.2
+ Medical Knowledge Initialization 76.0
+ Moderator’s Review 77.6
+ Moderator’s Review & MedRAG 80.3
Table 7: Impact of knowledge enhancement on MDAgents performance
We investigated whether simply assigning roles to agents is sufficient for expert-like performance, and explored the impact of equipping agents with different knowledge using Retrieval-Augmented Generation (RAG). Table 7 presents the results of these experiments.
These results indicate that while role assignment provides a foundation, augmenting agents with specific knowledge (using MedRAG) and structured reviews (Moderator’s Review) significantly enhances their ability to simulate domain expertise. The combination of Moderator’s Review and MedRAG yielded the best performance, highlighting the synergy between structured collaboration and domain-specific knowledge retrieval.
D.4 Complexity Assignment and Collaborative Settings
To address the impact of complexity assignment on accuracy and API costs, we conducted additional experiments focusing on high-complexity cases, particularly in image+text scenarios. Table 8 shows the results for various collaborative settings.
Collaboration Setting Accuracy (%)
Sequential & No Discussion 39.0
Sequential & Discussion 45.0
Parallel & No Discussion 56.0
Parallel & Discussion 59.0
Table 8: Impact of collaboration settings on high-complexity image+text tasks
These results underscore the importance of multi-turn discussions, particularly in complex cases. The parallel collaboration with discussion yielded the highest accuracy (59.0%), suggesting that enabling agents to work simultaneously and engage in dialogue is crucial for handling intricate medical queries. The significant performance gap between discussion and no-discussion scenarios (45.0% vs. 39.0% for sequential, and 59.0% vs. 56.0% for parallel) highlights the value of interactive deliberation in medical decision-making processes.